OpenStudy (anonymous):

If tan theta =12/5 and cos theta < 0, what is sin theta ? A. 12 b. 5/13 c. -15/13 d. 12/13 e.-12/13

OpenStudy (caozeyuan):

\[\tan \theta= \frac{ \sin \theta }{ \cos \theta } \]

OpenStudy (anonymous):

how do you find cos theta?

OpenStudy (caozeyuan):

\[\cos \theta <0 so \sin \theta <0\]

OpenStudy (caozeyuan):

no need to find cos theta, let me explain

OpenStudy (caozeyuan):

we now know that sin theta is less than 0 so we narrow the answers down to c or e

OpenStudy (caozeyuan):

but 15 cannot show up since the legs are 5,12 and the hypotenuse is 13

OpenStudy (caozeyuan):

so e is your answer! You see, you don't need to find cos theta

OpenStudy (caozeyuan):

simple elimination works, since it's a multiple choice question

OpenStudy (anonymous):

oh gotcha! i underestand. i just over confused myself. thank you :)

OpenStudy (anonymous):

can you help with one last one?